# Diffusion Beats Autoregressive in Data-Constrained Settings

### v2

@@ 3.2 Modeling Details for AR and Masked Diffusion
$\mathcal{L}_{\text{Diffusion}}\;=\;-\mathbb{E}_{r}\;\mathbb{E}_{\tilde{x}\sim q_{r}}\;\sum_{i\in\mathcal{M}}\log p_{\theta}\!\bigl{(}x_{i}\mid\tilde{x}\bigr{)}.$
Unlike Nie et al. (nie2024scaling, ), we do not add auxiliary losses (e.g., classifier-free guidance) in order to ablate only the factorization.
@@ 4.2 Fitting Data-Constrained Scaling Laws
Figure 3 illustrates how the utility of unique data decays with increased repetition. We evaluate this effect across three compute budgets—$1\times 10^{19}$, $3\times 10^{19}$, and $1\times 10^{20}$ FLOPs—by varying the proportion of unique data and parameters while keeping total compute fixed (e.g., 50% of the data for 2 epochs, 25% for 4 epochs, etc.). For each compute budget, we use single-epoch scaling laws to determine the optimal model size and unique token count for both AR and diffusion models. This experimental design allows us to directly measure how the utility of data diminishes with increased repetition. We present both empirical results and fitted curves from our parametric scaling law, observing strong agreement between the two. Notably, the decay rate of data value remains consistent across compute budgets for both model families. However, diffusion models consistently exhibit a substantially slower decay rate than AR models, suggesting they are better able to extract value from repeated data.
@@ 4.3 When to Use Diffusion over AR?
$\log_{10}(U)=0.460\cdot\log_{10}(C)-7.052,\quad\text{so}\quad C_{\text{crit}}(U)=2.12\times 10^{1.956}\cdot U^{2.174}.$
@@ 4.4 Downstream Results
Motivated by the critical compute threshold equation identified in Section 4.3, we scale the training data to 500M unique tokens and train a 2.3B parameter diffusion model using the compute budget predicted by the critical compute limit. The model was trained for 130 epochs, during which we observed no signs of convergence. Training was terminated due to compute constraints.
Across a diverse set of tasks and data scales, diffusion models consistently achieve higher accuracy than their AR counterparts. This validates our findings in Section 4.3, also confirms that the data efficiency gains observed in validation loss translate into stronger downstream performance. Additional results, including performance on textbook-style datasets, are provided in Table 3 in the Appendix.
@@ 5 Discussion
We hypothesize that the key advantage stems from the use of random factorization in diffusion models, which serves as a form of data augmentation. Unlike AR models, which are trained on a single, fixed left-to-right ordering, masked diffusion models are exposed to a wide variety of token prediction tasks and orderings during training—including, the left-to-right order itself. This broader distribution over prediction tasks encourages better generalization and more effective use of each training example. Analogous to image-based tasks where techniques like random cropping or color jittering boost generalization  caron2021emerging ; cubuk2020randaugment ; he2022masked , we believe the corruption process in diffusion plays a similar role. By learning to denoise across many corruptions, the model extracts richer signal per example over time, resulting in improved data efficiency.

### v3

@@ 3.2 Modeling Details for AR and Masked Diffusion
$\mathcal{L}_{\text{Diffusion}}\;=\;-\mathbb{E}_{r}\;\mathbb{E}_{\tilde{x}\sim q_{r}}\;\frac{1}{r}\sum_{i\in\mathcal{M}}\log p_{\theta}\!\bigl{(}x_{i}\mid\tilde{x}\bigr{)}.$
@@ 4.2 Fitting Data-Constrained Scaling Laws
[…]
@@ 4.3 When to Use Diffusion over AR?
$\log_{10}(U)=0.460\cdot\log_{10}(C)-1.050,\quad\text{so}\quad C_{\text{crit}}(U)=2.12\times 10^{1.956}\cdot U^{2.174}.$
@@ 4.4 Downstream Results
Motivated by the critical compute threshold equation identified in Section 4.3, we scale the training data to 500M unique tokens and train a 2.3B parameter diffusion model using the compute budget predicted by the critical compute limit. We train the model for 130 epochs, during which we observe no signs of convergence. We terminate the training due to compute constraints.
Across a diverse set of tasks and data scales, diffusion models consistently outperform their AR counterparts. This validates our findings in Section 4.3, also confirms that the data efficiency gains observed in validation loss translate into stronger downstream performance. Additional results, including performance on textbook-style datasets, are provided in Table 3 in the Appendix.
@@ 5 Discussion
We hypothesize that the key advantage stems from the random factorizations in diffusion objective, which serves as a form of implicit data augmentation. Unlike AR models, which are trained on a single, fixed left-to-right ordering, masked diffusion models are exposed to a wide variety of token prediction tasks and orderings during training—including, the left-to-right order itself. This broader distribution over prediction tasks encourages better generalization and more effective use of each training example. Analogous to image-based tasks where techniques like random cropping or color jittering boost generalization  caron2021emerging ; cubuk2020randaugment ; he2022masked , we believe the corruption process in diffusion plays a similar role. By learning to denoise across many corruptions, the model extracts richer signal per example over time, resulting in improved data efficiency. We further analyze this question in Section 9 in Appendix.